one machine in print shop can prodce a certain number of pages twice as fast as another machine.operating together,these machines can produce this number of pages in 8 mins.how long would it take each machine,working alone,to produce this number of pages?
Answer by checkley71(8403)

X*2X/(X+2X)=8
2X^2/3X=8 CROSS MULTIPLY
2X^2=24X
2X^2-24X=0
2X(X-12)=0
X-12=0
X=12 ANSWER.
2*12=24 ANSWER.
PROOF
12*24/3*12=8
288/36=8
8=8